A 5 kg block is sliding at an initial speed of 9 m/s across a surface, encountering a constant friction force of 6 N. How mucParts D and E only please
Answer #1

given
m = 5 kg
vi = 9 m/s
f = 6 N
d = 22 cm - 0.22 m

A) Workdone by friction, Wf = f*d*cos(180)

= 6*0.22*(-1)

= -1.32 J

B) Wnet = change in kinetic energy

Wnet = (1/2)*m*vf^2 - (1/2)*m*vi^2

2*Wnet/m = vf^2 - vi^2

==> vf = sqrt(vi^2 + 2*Wnet/m)

= sqrt(9^2 + 2*(-1.32)/5)

= 8.97 m/s

C) let x is the total distance tarvelled.

Wnet = KEf - KEi

f*x*cos(180) = 0 - KEi

-f*x = - (1/2)*m*v^2

x = (1/2)*m*v^2/f

= (1/2)*5*9^2/6

= 33.75 m

D) acceleration of the block during slowdown, a = -f/m

= -6/5

= -1.2 m/s^2

time taken to slowdown, t = (vf - vi)/a

= (0 - 9)/(-1.2)

= 7.5 s

Average power of friction = Workdone/time taken

= 6*33.75/7.5

= 27 W

E) instantaneous power = f*v

= 6*8.97

= 53.82 W

Note : Including sign the answers are

D) -27 W
E = -53.82 W